DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, October 6, 2004 Padelford C-36 3:50 pm Dirichlet-to-Neumann Map for Poincaré-Einstein Metrics

Robin Graham (UW)

The recent regularity theorem of Chruściel, Delay, Lee, and Skinner allows one to define an analogue of a Dirichlet to Neumann map for Poincaré-Einstein metrics (also known as asymptotically hyperbolic or conformally compact Einstein metrics). This talk will report on work in progress in understanding this map in even interior dimensions near the sphere, including explicit identification of its linearization at the sphere and consequences thereof.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.